OpenStudy (anonymous):

Let g(x) = x^2 + 2x - 6. Find the value of 2[g(2)]. Show your work.

OpenStudy (hoblos):

2[g(2)] = 2[ 2^2 + 2(2) -6 ] = 2[ 4+4 -6 ] = 2(2) = 4